OpenStudy (anonymous):

proportion- solve x-1/x+1=2/3x... i know you cross multiply to get (x-1)(3x)=(x+1)(2)...but then get stuck

OpenStudy (anonymous):

That's a great start. Use the distributive property to start simplifying each side of the equation you have. a(b + c) = ab + ac I'll do the left side of the equation for you. (3x)(x - 1) = (3x)(x) + (3x)(-1) = 3x^2 - 3x Now do the same for the right hand side and then solve for x.

OpenStudy (anonymous):

i'm confused..

OpenStudy (anonymous):

dont you cross multiply then factor out or something?

OpenStudy (anonymous):

No factoring. What mkuehn has shown you is the left side distributed. The same needs to be done on the right.

OpenStudy (anonymous):

So the right would be: \[2(x+1)=2x+2\]

OpenStudy (anonymous):

Then you bring evrything to one side to solve for 0.

OpenStudy (anonymous):

So your setup for the final solution would be: \[3x ^{2}-3x=2x+2\]

OpenStudy (anonymous):

i somehow got a squared x... i did 3x(x-1)=2(x+1) which brought me to 3x^2-3=2x+1

OpenStudy (anonymous):

or did i make that up

OpenStudy (anonymous):

You almost have it.

OpenStudy (anonymous):

3x * x = 3x^2 3x * -1 = ?

OpenStudy (anonymous):

so what i have now is 3x^2-3x=2x+2.. had some errors on my part

OpenStudy (anonymous):

Okay, so now bring the terms from the right hand side to the left hand side by subtracting them.

OpenStudy (anonymous):

then i add 3x to both sides?

OpenStudy (anonymous):

Yes, that's where I dropped off earlier. When doing on paper. I like to draw a line from the itkem being distributed to each item that needs to be multiplied so I can keep it straight.

OpenStudy (anonymous):

You can move everything to whichever side makes it easier on you, but they all need to be on one side.

OpenStudy (anonymous):

You can add 3x to both sides, but I think it is easier to work with the term that is squared as a positive, so bring everything to the left hand side.

OpenStudy (anonymous):

So if you add 3x to each side you also need to subtract 3x^2 as well.

OpenStudy (anonymous):

I agree with mkuehn on that. I prefer positive x^2 numbers as well :)

OpenStudy (anonymous):

Once you have them all on the same side you can factor if possible.

OpenStudy (anonymous):

Looking at this one I don't see an easy factoring so you're probably going to have to apply the quadratic formula.

OpenStudy (anonymous):

3x^2-5x-2=0?

OpenStudy (anonymous):

That's what I have as well. So ax^2+bx+c=0 Have you worked with the quadratic formula before?

OpenStudy (anonymous):

long time ago

OpenStudy (anonymous):

So we have: \[(-b \pm \sqrt{b ^{2}-4ac})/2a\]

OpenStudy (anonymous):

So we look at our current work. Can you find a,b,c?

OpenStudy (anonymous):

b=3 a=5 c=-2?

OpenStudy (anonymous):

Not quite :) The form is: \[ax ^{2}+bx+c=0\]

OpenStudy (anonymous):

ax^2+bx+c=0 3x^2-5x-2=0

OpenStudy (anonymous):

Go for it again.

OpenStudy (anonymous):

Keep in mind the signs matter. So a -5 will be a -5 in the quadratic.

OpenStudy (anonymous):

You're about 2 steps from your answers :)

OpenStudy (anonymous):

okay, do i add the -5x and -2?

OpenStudy (anonymous):

You need to get the values of a,b and c and then plug them into teh quadratic equation listed above. I'm trying to step you through getting a,b and c first. Then we'll plug them in.

OpenStudy (anonymous):

See how neatly they line up on mkuehn's last post. He's made it easier for you.

OpenStudy (anonymous):

If you give up I can give you a,b and c. I'm not trying to torture you. Just trying to heklp you through the process.

OpenStudy (anonymous):

yeah i understand that im just getting more confused as we go on....i'm trying to hard =[

OpenStudy (anonymous):

Okay looking at mkuehns post, a = 3 b = -5 c= -2

OpenStudy (anonymous):

Can you see that looking at it again?

OpenStudy (anonymous):

You've honestly done the hardest part earlier putting everything all on the same side :)

OpenStudy (anonymous):

Lots of folks get lost of the cross multiplying.

OpenStudy (anonymous):

So now we can go back to my earlier part and plug the numbers in: \[(-(-5)\pm \sqrt{(-5)^{2}-4(3)(-2)})/2(3)\]

OpenStudy (anonymous):

so -5+or- -5^2-4(3*-2)/2{3}?

OpenStudy (anonymous):

Simplifying we get: \[(5\pm \sqrt{25+24})/6\]

OpenStudy (anonymous):

Yes. Since we have x^2 in the problem there will be two answers. Thus the plus or minus part.

OpenStudy (anonymous):

Note how the -(-5) became 5. And the (-5)^2 became 25.

OpenStudy (anonymous):

That's why I mentioned the signs mattering.

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

So one more step and we'll have the two final answers: \[(5\pm \sqrt{49})/6\] So our answers are: \[(5+\sqrt{29})/6\] And \[(5-\sqrt{49})/6\] Since \[\sqrt{49}=7\] we can simplify some.

OpenStudy (anonymous):

That root 29 is supposed to be 49 I typoed

OpenStudy (anonymous):

\[(5+7)/6 = 12/6 = 2\]

OpenStudy (anonymous):

\[(5-7)/6 = -2/6 = -1/3\]

OpenStudy (anonymous):

So clear as mud? :P As you can see there was no way we could have factored out those answers.

OpenStudy (anonymous):

i dont get the last 2 posts

OpenStudy (anonymous):

\[\sqrt{49}\]

OpenStudy (anonymous):

=7

OpenStudy (anonymous):

So if you step up a set and ignore my 29 typo we replaced root 49 with 7 for both of our answers.

OpenStudy (anonymous):

The easy way to test is to plug the answers back into: \[3x ^{2}-5x-2=0\] and make sure you get 0. So the first answer was 2: \[3(2)^{2}-5(2)-2=3(4)-10-2=12-12=0\]

OpenStudy (anonymous):

The second answer was -1/3: \[3(-1/3)^{2}-5(-1/3)-2=3(1/9)+5/3-2=3/9+5/3-2=1/3+5/3-2=6/3-2=2-2=0\]

OpenStudy (anonymous):

That last part that got cut off is: \[6/3-2=2-2=0\]

OpenStudy (anonymous):

i dont even know what i'm looking at

OpenStudy (anonymous):

Okay we solved using the quadratic equation with a, b, and c that you helped find.

OpenStudy (anonymous):

We plugged them into that formula with the plus and minus which we separated into two parts.

OpenStudy (anonymous):

You're looking at magi... err math. :)

OpenStudy (anonymous):

When we solved those parts we found the answers to be 2 and -1/3.

OpenStudy (anonymous):

From there we wanted to make sure we found the right answers and didn't make a mistake so we plugged it back into the original formula.

OpenStudy (anonymous):

using the formula

OpenStudy (anonymous):

That what the last part was. I know it all looks intimindating, but so does a multicourse meal. Small bites and you'll get there :)

OpenStudy (anonymous):

I have to roll. I know it seems a little intimidating but it's not as bad as it looks. Doing a few more of the problems may make it make more sense. When I come across math problems that leave me lost I head over to khanacademy.org. You may want to give it a look see. Sometimes things make more sense when you see them visually solved.

OpenStudy (anonymous):

yeah i'll try that. thank you very much though for spending so much time on that problem